OpenStudy (anonymous):

Factor the polynomial 8x^4+6x^3-2x^2 @dmezzullo @Destinymasha @waterineyes @SnuggieLad @So1D @smokeydabear @sar12 @primeralph @@erikgomez @Zale101 @Zarkon

OpenStudy (anonymous):

Nice way of tagging all in the question itself.. :)

OpenStudy (anonymous):

1. Firstly you can factor out \(x^2\) out of it..

OpenStudy (anonymous):

\[\implies \color{green}{x^2(8x^2 + 6x - 2)}\] Now you have got Quadratic equation, can you factor this equation or not?

OpenStudy (anonymous):

2x^2(x+1)(4x-1)

OpenStudy (anonymous):

Then I would bring out the 2 from waterineyes's equation.

OpenStudy (anonymous):

Yes you can..!!

OpenStudy (primeralph):

yeah, yeah.........since ya'll got it, can I troll?

OpenStudy (anonymous):

Go for it, the floor is yours

OpenStudy (anonymous):

8x^4+6x^3-2x^2 =2x^2(4x^2+3x-1) =2x^2[4x^2+4x-1x-1] (4*-1=-4,4-1=3) now you can factorise

OpenStudy (anonymous):

Someone has factorized it earlier, so no need to again factorize it.. :)

OpenStudy (primeralph):